# Fine-Tuning Universal Machine-Learned Interatomic Potentials: A Tutorial on Methods and Applications

### v1

@@ III.2 Molybdenum
III.3 Graphene–Water Interface
@@
Figure 8: Comparison of graphene–water interface simulations using MACE-FT-MD and AIMD. (a) Temperature evolution over time. (b) Oxygen number-density profile along the z-axis; The first peak corresponds to the graphene–water distance, ${\rm d}_{\rm gw}=0.35~{}{\rm nm}$, which is in good agreement with the experimental value of 0.36 nm [109]. The inset illustrates the definition of ${\rm d}_{\rm gw}$.
Table 7 reports the computational cost (wall-clock time in seconds) for 10,000-step molecular dynamics simulations performed with and without the cuEquivalence kernel, across systems of varying sizes (128 to 2900 atoms). For each configuration, simulations using cuEquivalence (“With”) show consistently lower elapsed times compared to runs without it (“Without”), especially for larger systems. For instance, at 2900 atoms, cuEquivalence reduces the runtime from 5084 seconds to 2619 seconds. The scaling factor improves from 0.76 (without cuEquivalence) to 0.41 (with cuEquivalence), indicating better parallel efficiency with the kernel enabled.
@@ IV Summary and Perspective
This tutorial provides a practical guide to fine-tuning universal machine-learned interatomic potentials (U-MLIPs), using the MACE-MP-0 model as a representative example. We cover the end-to-end process—from dataset preparation and hyperparameter selection to model training and application—with the goal of helping researchers efficiently adapt pre-trained models to their systems of interest. The effectiveness of fine-tuning is demonstrated through representative examples, including solid-state electrolytes, stacking fault defects in metals, and surface interactions in low-dimensional materials. To support reproducibility, all code and data are made publicly available, along with scripts for running molecular dynamics simulations using the RBMD platform.
@@
Long-Range Interactions.
Most MLIPs rely on a local decomposition of the total energy, in which each atom’s contribution is determined by its local environment within a finite cutoff radius (typically 5–10 Å). While this approximation is effective in condensed-phase systems where long-range interactions are screened, it becomes inadequate in systems such as ionic solids, electrolytes, and interfacial structures, where electrostatic and dispersion interactions extend well beyond the cutoff. In these cases, long-range effects may be misinterpreted as noise during training, ultimately degrading model accuracy. To address this issue, several extensions have been proposed [110, 111, 112, 113, 114], including environment-dependent charge models, DFT-D-type dispersion corrections, and hybrid approaches that integrate ML models with explicit physics-based components. Continued progress in this direction is expected, as long-range interactions remain a key challenge and an active area of research in the development of accurate and transferable MLIPs.
Multiple Electronic States.
Standard MLIPs are designed to model a single potential energy surface, typically the ground state under the Born–Oppenheimer approximation [115]. However, many important phenomena—including spin transitions, charge transfer, and photocatalysis—require access to excited-state dynamics. Representing multiple electronic states remains a challenge due to the need to model sharp energy differences and nonadiabatic couplings. Some approaches train separate ML models for different states [116, 117, 118], while others explore multi-output architectures [119]. Recent work [120] has also focused on spin-dependent descriptors and graph neural networks with embedded electronic information, which show promise in this direction.
@@
After pre-training a foundation model, post-training serves as a crucial step to further adapt and optimize the model for specific downstream applications. One effective post-training technique is model distillation, in which a simpler model (the “student”) is trained to replicate the behavior and predictive performance of a larger, more complex model (the “teacher”—in this case, U-MLIPs). This is typically achieved by using the teacher’s outputs or internal representations as supervisory signals for training the student [125]. Recent work [126] demonstrates that matching the Hessians of energy predictions between the teacher and student enables effective distillation. In particular, large-scale foundation models can be distilled into compact, task-specific MLIPs tailored to a particular chemical subdomain, achieving up to 50$\times$ speedup in inference time compared to the original foundation model. Future investigations on this topic will be conducted using the RBMD platform.
V Appendices
This appendix provides supplementary materials to complement the main content of the tutorial. Section V.1 presents a concise overview of existing U-MLIPs, offering historical context and highlighting key differences among representative models. Section V.2 provides additional technical details on the MACE architecture for readers seeking a deeper understanding of its design and capabilities. Finally, Section V.3 compiles a list of useful code repositories and publicly available datasets to support reproducibility and facilitate the practical application of the methods described in this work.
V.1 Review of U-MLIPs
While recent advances in universal machine-learned interatomic potentials (U-MLIPs) are largely driven by modern machine learning techniques, their conceptual roots can be traced to earlier semi-universal, physics-based force fields (SU-PBPs). Early examples of SU-PBPs include the work of Weiner et al. [127] in 1981, which laid the foundation for general-purpose molecular potentials. This line of development continued with the introduction of widely used models such as the Universal Force Field (UFF) by Rappé et al. [128] and the AMBER family of force fields [127, 129], both of which have seen broad application in molecular simulations, particularly in fields like pharmaceutical research.
Despite their shared aim of general applicability, SU-PBPs and U-MLIPs differ significantly in methodology and maturity. While SU-PBPs are based on empirical and physically motivated functional forms, U-MLIPs are trained directly on quantum mechanical data using flexible machine-learned architectures. A pioneering example of a semi-universal MLIP is the ANI (Accurate NeurAl network engINe for Molecular Energies) model, introduced by Smith et al.[11, 130]. Originally trained on organic molecules containing C, H, O, and N, ANI was later expanded to incorporate S, Cl, and F, thereby achieving coverage of approximately 90% of drug-like chemical space[131]. Functionally, ANI can be seen as a machine-learned counterpart to traditional force fields like AMBER.
Further developments have extended the scope and accuracy of ANI-type models. For instance, AIMNet2 [132] includes 14 elements and introduces corrections for long-range interactions, such as dispersion and electrostatics, enabling improved modeling of charged systems and diverse bonding environments. Similarly, ANI-1xnr [133] adapts the ANI framework for use in condensed-phase systems, including liquids and reactive environments.
More recently, the MACE-OFF23 model [134] has leveraged the MACE message-passing architecture to construct a short-range, transferable potential for the ten most prevalent elements in organic chemistry (H, C, N, O, F, P, S, Cl, Br, I). Despite not explicitly modeling long-range interactions, MACE-OFF23 has shown performance improvements over ANI across several benchmarks, highlighting the progress in developing efficient and accurate ML-based universal potentials.
The first truly universal MLIPs with broad elemental coverage across the periodic table emerged around 2022 [27, 135]. Since then, many U-MLIPs have been proposed, including:
  • M3GNet [27]: a 3-body materials graph network developed by Chen and Ong;
  • CHGNet [23]: the Crystal Hamiltonian Graph Neural Network;
  • ALIGNN-FF [26]: a unified atomistic force field based on line graph neural networks;
  • PFP [33]: a tensorial message-passing network developed by Takamoto et al., also released commercially in Matlantis;
  • GNoME [24]: based on NequIP [8], trained on $\sim$80 million in-house DFT calculations;
  • SevenNet-0 [34]: optimized for large-scale simulation with better parallel scaling;
  • MACE-MP-0 [6]: a U-MLIP based on the MACE architecture [6], trained on the same public data as CHGNet and SevenNet, demonstrating strong performance across aqueous systems, catalysis, and MOFs;
  • GPTFF [136]: a GNN-transformer hybrid force field;
  • MatterSim [35]: trained by Microsoft on a large dataset of $\sim$17 million structures covering broad thermodynamic conditions;
  • Orb [137]: a high-speed CUDA-optimized model offering 3–6$\times$ speedup over PyTorch-based MACE;
  • Equiformer V2-OMAT24 [36]: currently ranked highest on the MatBench leaderboard, trained on $\sim$118 million structures.
There are several recent studies have systematically benchmarked U-MLIPs [138, 139, 41, 140]. These evaluations underscore the importance of assessing model performance not only in terms of energy and force errors, but also with respect to physically meaningful properties such as surface energies, elastic moduli, and defect energetics—quantities that are often more relevant for practical applications. However, constructing diverse and high-quality test sets for such evaluations remains an open challenge. A key concern is the risk of overfitting when test datasets become implicit targets during model optimization, which may compromise generalization to unrelated properties or systems. Moreover, it is essential to recognize that low errors in energy and force predictions do not necessarily ensure reliable performance on derived physical properties. This limitation is inherent to MLIPs in general and is not specific to U-MLIPs.
V.2 MACE Architecture
@@
V.3 Codes and Data Availability
To support reproducibility, extensibility, and practical adoption of MACE-based workflows, Table 8 summarizes key publicly available GitHub repositories and data resources. These include official codebases for model construction and training, fine-tuning utilities, simulation interfaces (e.g., LAMMPS [16], ASE [15], OpenMM [144]), and model conversion or acceleration tools. The collection also contains pretrained foundation models and benchmark datasets. Together, these components provide a comprehensive ecosystem for developing and deploying MACE across diverse material systems.
Table 8: Key repositories and data resources supporting the development, deployment, and application of MACE-based models. The table includes core codebases, pretrained models, tutorials, simulation interfaces, kernel and dispersion extensions, and benchmark datasets.
Repository / Resource Purpose / Description Link / Owner
mace Official MACE codebase for model construction, training, evaluation, and inference. ACEsuit/mace
MACE-MP-0 Pretrained MACE-MP-0 foundation model checkpoints, logs, and metadata. ACEsuit/mace-foundations
mace-off MACE-OFF23 pretrained transferable organic force fields for molecular systems. ACEsuit/mace-off
mace-ft-tutorial Fine-tuning tutorial with scripts and examples for adapting pretrained MACE models. https://github.com/John2021-hub/mace-ft-tutorial
mace-tutorials-csc A tutorial for training MACE models. ilyes319/mace-tutorials-csc
mace-jax JAX-based reimplementation of MACE for TPU/GPU acceleration and efficient autodiff. ACEsuit/mace-jax
mace-layer MACE layer for neural operator applications on 3D geometric point clouds. ACEsuit/mace-layer
mace-lammps LAMMPS interface for deploying MACE potentials in molecular dynamics simulations. ACEsuit/mace-lammps
mace-md MD wrapper supporting MACE potentials, especially with OpenMM backend. jharrymoore/mace-md
RBMD Random-batch MD framework for large-scale atomistic simulations using MACE. https://www.randbatch.com/rbmd
torch-dftd PyTorch-based DFT-D3 module for dispersion correction, usable via ASE with MACE. CheukHinHoJerry/torch-dftd
cuEquivariance CUDA-accelerated kernel library for equivariant tensor operations in MACE. NVIDIA/cuEquivariance
cuda-mace Custom CUDA implementation of MACE for high-performance inference. nickjbrowning/cuda_mace
Materials Project (MP) data Dataset used to train MACE-MP-0 and CHGNet; includes DFT-calculated material properties. materialsproject.org

### v2

@@ III.3 Molybdenum
III.4 Graphene–Water Interface
@@
The graphene–water dataset used to fine-tune the U-MLIP was generated by pre-equilibrating a periodic graphene–water interface with classical molecular dynamics (LAMMPS) [113, 114]. Graphene was modeled with a Tersoff bond-order potential [115, 116], and water with a rigid three-site model constrained by SHAKE; short-range interactions were described by Lennard–Jones terms [117], and long-range electrostatics by Ewald summation under periodic boundary conditions. Canonical (NVT) sampling furnished equilibrated configurations, from which temporally uncorrelated snapshots were randomly selected for ab initio labeling. Single-point DFT labels were obtained in VASP [118] using the PAW [119] formalism and the PBE exchange–correlation functional [120] augmented with D3 dispersion [121] to capture graphene–water physisorption. cutoff of 400 eV was chosen for the plane‑wave basis, and electronic self‑consistency was achieved when the total‐energy change fell below $1\times 10^{-6}$ eV. Geometry optimisations were considered converged once the residual forces on all atoms were smaller than $0.02$ eV Å-1.
@@
Figure 9: Comparison of graphene–water interface simulations using MACE-FT-MD, MACE-MPA-0-MD and AIMD. (a) Temperature evolution over time. (b) Oxygen number-density profile along the z-axis; The first peak corresponds to the graphene–water distance, ${\rm d}_{\rm gw}=0.35~{\rm nm}$, which is in good agreement with the experimental value of 0.36 nm [122]. The inset illustrates the definition of ${\rm d}_{\rm gw}$. To make a comparison, we also plot the MD trajectory for MACE-MPA-0 model without fine-tuning.
Table 7 reports the computational cost (wall-clock time in seconds) for 10,000-step molecular dynamics simulations performed with and without the cuEquivalence kernel, across systems of varying sizes (128 to 2900 atoms). For each configuration, simulations using cuEquivalence (“With”) show consistently lower elapsed times compared to runs without it (“Without”), especially for larger systems. For instance, at 2900 atoms, cuEquivalence reduces the runtime from 5084 seconds to 2619 seconds. The scaling factor improves from 0.76 (without cuEquivalence) to 0.41 (with cuEquivalence), indicating better parallel efficiency with the kernel enabled. These results suggest that the new kernel can serve as an efficient building block for large-scale and long-timescale molecular dynamics simulations.
@@
III.5 Molecule/Oxide Interface
As a representative case of molecule–oxide interactions, we consider a gold dimer (Au2) adsorbed on the MgO(001) surface with aluminum doping, following [123]. The dataset is also from the authors of  [123]. This system has been widely employed to benchmark long-range effects such as charge transfer and polarization at oxide surfaces. It thus provides a stringent test of whether fine-tuned foundation models can accurately capture long-range physics beyond the reach of conventional MLIPs.
Figure 10 compares the performance of the foundation MACE model (MACE-MPA-0) with its fine-tuned counterpart (MACE-FT): (a) energy RMSE per atom (meV/atom), and (b) force RMSE (meV/Å). In both metrics, MACE-FT achieves substantially lower errors than MACE-MPA-0, demonstrating improved agreement with the reference data (lower is better). The relatively large error of MACE-MPA-0 arises because its reported value corresponds to the step-0 output during training, reflecting a mismatch between the DFT settings used in this dataset and those employed in the original MPA training. Notably, even without explicitly modeling long-range interactions, the fine-tuned foundation model outperforms recent state-of-the-art methods reported in [123]. The CACE-LR model, which incorporates a latent Ewald summation to capture long-range effects, already delivers strong accuracy; yet our fine-tuned model achieves comparable or even superior performance. This suggests that fine-tuning can implicitly adapt to capture aspects of long-range physics, highlighting the potential of foundation models as a flexible and powerful basis for tackling complex interfacial systems.
Figure 10: Performance comparison between a foundation MACE model (MACE-MPA-0) and its fine-tuned counterpart (MACE-FT) on the doped Au-MgO surface. (a) Atomic structure; (b) energy RMSE per atom (meV/atom); (c) force RMSE (meV/Å). MACE-FT achieves consistently lower errors than MACE-MPA-0 and reaches accuracy comparable to CACE-LR [123], a state-of-the-art method with explicit long-range interactions.
III.6 Solid–Solid Interface
Atomistic modeling of solid–solid interfaces is essential for understanding the synthesizability and stability of materials [124]. Such interfaces are notoriously challenging to model because their heterogeneous nature often requires long-period supercells and an accurate treatment of long-range interactions, including charge transfer and polarization, which are typically beyond the scope of conventional MLIPs. To assess whether fine-tuned foundation models can capture these effects, we consider a representative heterointerface from [123], namely LiCl(001)/GaF3(001). The atomic structure is showin in Figure 11.
Figure 11: Atomic structure of the LiCl(001)/GaF3(001) interface used in this study.
Table 8 reports the test RMSEs for energies and forces before and after fine-tuning, together with reference results from [123]. The CACE-LR method, which employs a latent Ewald summation to explicitly describe long-range interactions, already achieves impressive accuracy. Remarkably, however, the fine-tuned foundation model (MACE-FT) performs even better, despite lacking any explicit long-range correction. We attribute this to the fact that large pre-trained models may implicitly encode aspects of long-range physics during pre-training, which fine-tuning then leverages and refines for the target interface. This observation highlights an intriguing direction: foundation models, when fine-tuned, might generalize long-range effects in ways not captured by traditional MLIPs. We plan to further investigate this phenomenon in future work.
Table 8: Test RMSEs for energies (E, meV/atom) and forces (F, meV/Å) of the LiCl(001)/GaF3(001) interface. Results for CACE-LR are taken from [123]; “–” indicates values not reported. The fine-tuned MACE-MPA-0 (MACE-FT) achieves the best accuracy across both metrics.
MACE-MPA-0 CACE-LR MACE-FT
E (meV/atom) 97.04 0.09
F (meV/Å) 179.2 67.8 18.23
@@ IV SUMMARY AND PERSPECTIVE
This tutorial provides a practical guide to fine-tuning universal machine-learned interatomic potentials (U-MLIPs), using the MACE-MP-0 model as a representative example. We cover the end-to-end process—from dataset preparation and hyperparameter selection to model training and application—with the goal of helping researchers efficiently adapt pre-trained models to their systems of interest. The effectiveness of fine-tuning is demonstrated through representative examples, including solid-state electrolytes, stacking fault defects in metals, surface interactions in low-dimensional materials, and more complicated heterointerfaces. To support reproducibility, all code and data are made publicly available, along with scripts for running molecular dynamics simulations using the RBMD platform.
@@
[…]
V Acknowledgements
We gratefully acknowledge the original developers of MACE, particularly Prof. Gábor Csányi’s group, for their insightful discussions and valuable contributions. We also appreciate the helpful suggestions and exchanges on the MACE GitHub repository (https://github.com/ACEsuit/mace) that have informed and improved this tutorial on fine-tuning MACE foundation models. Readers are encouraged to consult the official MACE documentation (https://mace-docs.readthedocs.io/en/latest/), upon which parts of this tutorial are based. Finally, we gratefully acknowledge the author of the GitHub repository (https://github.com/BingqingCheng/cace-lr-fit) for generously sharing the dataset that enabled our last two numerical examples.
VI Codes and Data Availability
The datasets and source code supporting all numerical results presented in this work are publicly available in the tutorial repository: https://github.com/John2021-hub/mace-ft-tutorial.git.
VII MACE Architecture